Although gas station owners lock their tanks at ni
Answer #1

a) The lowest pressure that a vacuum pump could ever hope to attain is 0 bar
The height of a column of fluid (assuming the vapour pressure in the tank is 1 bar) then maximally is
Delta(p) = rho g Delta(h)
Delta(h) = Delta(p) / (rho g)
= 10^5 N/m^2/(733 kg/m^3 * 9.81 m/s^2) = 13.9 meter = 45.6 ft


b) If Delta(p) is not 1 bar but ( 1 - 0.194) bar = 0.806 bar, then the depth needed is reduced by a factor 0.806

So then
D = 0.806 * 45.6 feet = 36.75 feet